# Effect of noise characterization on the detection of mHz stochastic gravitational waves

### v1

@@ 4.1 Impact of prior and model choice
In figure 2 we plot the Bayes Factors $B_{10}$ between the noise-plus-signal $H_{1}$ and the noise-only $H_{0}$ hypotheses for injected power-law signals of varying amplitudes and spectral indices. Each row corresponds to a different analysis model. The first row shows the results for the full flexible non-parametric models for both signal and noise. In the second row, we show the results using a template for the signal, while the noise description remains non-parametric. In the third row, we present the results of the analyses using the analytic models for signal based on Eq. (3.11) and noise based on Eqs. (3.9) and (3.10).
Each column corresponds to a different noise prior. The left-hand side column shows results with the informative (skewed-normal) prior, while the right-hand side column shows results with the non-informative (uniform) prior. For all analyses, we have chosen the conservative limit of $B_{10}=100$ as the threshold for a confident detection of a stochastic signal [23].
@@
Although the choice of model type for the signal (parametric and non-parametric) brings an improvement, it has a smaller effect on the detectability bounds on the parameter space than the choice of prior. In figure 3a we present a direct comparison of the impact of the different models used in the analysis, using different line styles. As expected, a data-consistent parametric template for the signal enables the detection of a SGWB in a larger portion of the parameter space when comparing to the non-parametric case. This can be attributed to the lower dimensionality of the model, which simultaneously controls and constrains the different stochastic contributions across the frequency range. When using a parametric (analytic) model for both the noise and the signal, this advantage diminishes, as shown by the overlapping dashed-dotted lines of figure 3a. In figure 3b, we show the ratio between the detectability bounds assuming different priors for the PSD of the noise. It is the ratio between the solid lines of different colors of figure 3a. Examples of the reconstruction of the injected stochastic signal are given in figure 4, while more information can be found in the Appendix A, and in its figures A1, where we compare the parameter estimates of the SGWB signal with a power-law model under different prior assumptions. In figure A2, we show an example of the marginal posterior of the spline knots amplitude and position on the frequency axis for the two noise components of the instrument.
Figure 3: (a) The detectability limits depending on the adopted type of models for the noise signal and the prior used on the noise components. The color denotes the type of prior used for the noise PSD, while the line style denotes the type of model. This panel is basically a summary of figure 2. (b) Comparison of the detectability limits for different types of models with respect to the assumed prior on the spectral shape of the noise. This figure essentially presents the ratio between the curves of the two column panels of figure 2, or the ones from the left panel of this figure. The choice of informative prior is more impactful at the lower-part of the frequency spectrum, where the noise and signal transfer functions are comparable. We note similar behavior for both types of model for the SGWB. This does not hold true for the case where analytic models where adopted for all components of the data covariance matrix (orange line).
(a) (Left): Non-parametric models for both the signal and the noise, with a more informative prior on the noise spectral shape, based on the skewed-normal distribution (see 3.6 for details). (Right): Same as in left panel, but after adopting a broad uniform prior on the noise spectral shape.
(b) (Left): Non-parametric model for the noise, and a parametric model for the SGWB signal. The informative skewed-normal prior on the noise spectral shape was used. (Right): Same as in left panel, but after adopting a broad uniform prior on the noise spectral shape.
(c) (Left): Parametric models for both the SGWB signal and the noise, and skewed normal prior on the spectral shape of the noise. (Right): Same as in left panel, but after adopting a broad uniform prior on the noise spectral shape.
Figure 4: Examples of signal reconstruction for an injection of a SGWB that follows the shape of a power-law with with $\Omega_{0}=5\times 10^{-15},\,n=2$. The analysis has been performed under different assumptions about the noise priors and SGWB spectral models.
@@ 5 Discussion
The first type of models used is the parametric (with $\theta_{\mathrm{GW}}\equiv\{\Omega_{0},\,n\}$ for the signal and $\theta_{\mathrm{n}}\equiv\{a_{\mathrm{TM}},\,a_{\mathrm{OMS}}\}$ for the noise), for which their spectral shape is considered known apriori. The second type is the more flexible non-parametric model based on spline interpolation methods. This type of model is able to capture unmodeled features on the spectral shape of the data, allowing for the search and reconstruction of virtually any SGWB signal or instrumental noise. The optimal model order is inferred from the data with the use of trans-dimensional sampling methods (RJ-MCMC and the Eryn package [31]). In both cases, we model the different components of the data covariance matrix at the link level, and then we apply the known TDI transformation matrices. This safeguards against spurious effects caused by the imperfect modelling of the given spectrum around the areas where it falls to zero. At the same time, we avoid modelling complicated features of the spectra, which directly translates to simpler models with smaller dimensionality.
@@
As expected, an informative prior knowledge of the instrumental noise greatly affects the detectability of the underlying SGWB signal, allowing for a greater portion of the parameter space to be measured by LISA. Knowing the spectral shape of the noise brings an improvement, as evident from the lower two panels of figure 2. A parametric model of the noise allows for improved resolvability of the underlying signal, as opposed to the more flexible non-parametric models based on spline interpolation methods. On the other hand, the choice of model type of the SGWB signal (parametric vs. non-parametric) does not have as much impact as the prior choice. The above results are expected to vary under more realistic conditions, e.g., with unequal noises across the different spacecraft of the constellation. Similarly, a time-frequency analysis must also be adopted for more realistic orbits. In this work, we use simplified analytic Keplerian orbits, which allow us to assume a constant transfer function for the duration of the measurement across the frequency range.
Finally, we have assumed conservative uncertainties in the spectral estimates, as we adopt a lower bound for the effective number of degrees of freedom in the Wishart likelihood. This can be refined in the future by accurately including the effect of windowing.
The pipeline presented in this work is versatile and can fit various types of models while adapting their dimensionality. In addition, it can also handle time-varying stochastic components, making it suitable for inferring both the non-stationary properties of the instrumental noise as well as anisotropic SGWB signals. A prime example is the stochastic signal generated by the unresolved Compact Galactic Binaries. We believe that such a tool will be useful for the analysis of the future data of the LISA mission.
@@ Acknowledgments
We acknowledge useful discussions with F. Pozzoli, R. Buscicchio, M. Muratore, N. Dam Quang, O. Hartwig, M pieroni, A. Santini, S. Babak, C. Caprini, H. Inchauspé, A. Petiteau, M. Besançon, and G. Nardini and the members of the LISA Distributed Data Processing Center (DDPC).
@@
For the analysis and modelling of the LISA constellation we have used the backgrounds software, which can be accessed via a dedicated gitlab repository:
https://gitlab.in2p3.fr/qbaghi/backgrounds. We have also utilized the CudAkima package by A. Santini [40], and the Eryn sampler for its RJ-MCMC capabilities.
@@ Appendix A Complementary figures for GW parameter posteriors
In this section we present complementary figures that are relevant to the analysis performed for the production of figure 2. As already mentioned in section 3, we perform parameter estimation analyses for different injections of power-law SGWB signals, under different prior assumptions on the instrumental noise, while using analytic models and flexible non-parametric models for the signal and noise components of the data. As an example, in figure A1 we show the parameter estimation results for a stochastic signal with $\Omega_{0}=5\times 10^{-15},\,n=2$. By the result shown in figure 2, we should expect this signal to be detectable with very high confidence when using analytic models for all components (red data of figure A1), while the detectability to drop when using the more spline models for the PSD of the noise components (blue data of same figure). The two panels of figure A1 demonstrate the effect of the prior choice.
@@
In figure A2 we show the resulting combined posteriors on the spline knots positions and amplitudes for each model order $k$, as sampled with trans-dimensional MCMC methods. The left panel corresponds to the TM, while the right panel to the OMS noise components. We remind the reader that all noises are set to be equal across the three spacecraft, and therefore are being fit simultaneously with the same model in our analysis. The shaded gray area represents the wide uniform prior around the injected signal. Since we fit for deviations from the injected spectra, there is no need for high-order models to capture the complexity of the data, and therefore the number of spline knots estimated converges to the lower prior bound of $k=2$. This is shown in the embedded histograms in both panels of Figure A2.
@@
Figure A1: Comparison of the parameter estimates of the SGWB signal ($\Omega_{0}=5\times 10^{-15},\,n=2$) with a power-law model, under different prior assumptions, and with different models for the noise (parametric and non-parametric, based on spline interpolation). (a) Using the informative prior based on the $\mathcal{SN}$ distribution, and (b) using the flat uninformative prior. See text for details.
@@
Figure A2: Examples of the combined posterior of the spline knots amplitudes $\log S_{j,k}$ and position $\log f_{j,k}$ for each $j$ spline number and given model order $k$, for the two instrumental noises. (a) for the TM and (b) the OMS noises. The shaded area represents the uniform prior used for this particular run, while the dashed black line denotes the injected true spectrum. The embedded figure represents the estimated model order, or the number of spline knots, as estimated via RJ-MCMC sampling. This analysis corresponds to a signal injection with $\Omega_{0}=5\times 10^{-15},\,n=2$ and the final signal estimate shown in figure A1b. See text for more details.

### v2

@@ 4.1 Impact of prior and model choice
Figure 3: (a) The detectability limits as functions of the adopted type of models for the noise signal and the prior used on the noise components. The color denotes the type of prior used for the noise PSD, while the line style denotes the type of model. This panel is basically a summary of figure 2. (b) Comparison of the detectability limits for different types of models with respect to the assumed prior on the spectral shape of the noise. This figure presents the ratio between the curves of the two column panels of figure 2, or the ones from the left panel of this figure. The choice of informative prior is more impactful at the lower part of the frequency spectrum, where the noise and signal transfer functions are comparable. We note a similar behavior for both types of SGWB model. This does not hold true for the case where analytic models where adopted for all components of the data covariance matrix (orange line).
In figure 2 we plot the Bayes Factors $\mathcal{B}_{10}$ between the noise-plus-signal $H_{1}$ and the noise-only $H_{0}$ hypotheses for injected power-law signals of varying amplitudes and spectral indices. Each row corresponds to a different analysis model. The first row shows the results for the full flexible models for both signal and noise. In the second row, we show the results using a template for the signal, while the noise description remains flexible. In the third row, we present the results of the analyses using the analytic models for signal based on eq. (3.11) and noise based on eqs. (3.9) and (3.10).
Each column corresponds to a different noise prior. The left-hand side column shows results with the informative (skewed-normal) prior, while the right-hand side column shows results with the non-informative (uniform) prior. For all analyses, we have chosen the conservative limit of $\mathcal{B}_{10}=100$ as the threshold for a confident detection of a stochastic signal [25].
(a) (Left): Flexible models for both the signal and the noise, with a skewed-normal prior on the noise spectral shape. (Right): Same as in left panel, but with a broad uniform prior on the noise.
(b) (Left): Flexible model for the noise, parametric model for the SGWB signal, using the skewed-normal prior on the noise. (Right): Same as in left panel, but with a uniform prior on the noise.
(c) (Left): Parametric models for both the SGWB signal and the noise, and skewed normal prior on the noise. (Right): Same as in left panel, but with a uniform prior on the noise.
Figure 4: Examples of signal reconstruction for an injection of a SGWB that follows the shape of a power-law with with $\Omega_{\mathrm{GW}}=5\times 10^{-15},\,n=2$. The analysis has been performed under different assumptions about the noise priors and SGWB spectral models.
@@
Although the choice of model type for the signal (parametric and flexible) brings an improvement, it has a smaller effect on the detectability bounds on the parameter space than the choice of prior. In figure 3a we present a direct comparison of the impact of the different models used in the analysis, using different line styles. As expected, a data-consistent parametric template for the signal enables the detection of a SGWB in a larger portion of the parameter space when comparing to the flexible case. This can be attributed to the lower dimensionality of the model, which simultaneously controls and constrains the different stochastic contributions across the frequency range. When using a parametric (analytic) model for both the noise and the signal, this advantage diminishes, as shown by the overlapping dashed-dotted lines of figure 3a. In figure 3b, we show the ratio between the detectability bounds assuming different priors for the PSD of the noise. It is the ratio between the solid lines of different colors of figure 3a. Examples of the reconstruction of the injected stochastic signal are given in figure 4, while more information can be found in the Appendix A, and in its figures A1, where we compare the parameter estimates of the SGWB signal with a power-law model under different prior assumptions. In figure A2, we show an example of the marginal posterior of the spline knots amplitude and position on the frequency axis for the two noise components of the instrument. Finally, in figure A3 we show an example where a marginally detectable isotropic stochastic signal can be misinterpreted as excessive power of the TM noise, especially when using flexible spectral models.
@@ 5 Discussion
The first type of models used is a fully parametric model (with $\boldsymbol{\theta}_{\mathrm{GW}}\equiv\{\Omega_{\mathrm{GW}},n\}$ for the signal and $\boldsymbol{\theta}_{\mathrm{n}}\equiv\{a_{\mathrm{TM}},a_{\mathrm{OMS}}\}$ for the instrument noise), for which the spectral shape is considered known apriori. The second type is the more flexible model based on spline interpolation methods. This type of model is able to capture unmodeled features on the spectral shape of the data, allowing for the search and reconstruction of virtually any SGWB signal or instrumental noise. The optimal model order is inferred from the data with the use of trans-dimensional sampling methods (RJ-MCMC and the Eryn package [35]). In both cases, we model the different components of the data covariance matrix at interferometer level, and apply the known transformation matrices to form the TDI combinations. This safeguards against spurious effects caused by the imperfect modelling of the TDI spectrum around the areas where it falls to zero. At the same time, we avoid modelling complicated features of the spectra, which directly translates to simpler models with smaller dimensionality.
@@
As expected, an informative prior knowledge of the instrumental noise greatly affects the detectability of the underlying SGWB signal, allowing for a greater portion of the parameter space to be measured by LISA. Knowing the spectral shape of the noise brings an improvement, as evident from the lower two panels of figure 2. A parametric model of the noise allows for improved resolvability of the underlying signal, as opposed to the more flexible models based on spline interpolation methods. On the other hand, the choice of model type of the SGWB signal (parametric vs. flexible) does not have as much impact as the prior choice. The above results are expected to vary under more realistic conditions, e.g., with unequal noises across the different spacecraft of the constellation. Similarly, a time-frequency analysis must also be adopted for more realistic orbits. Indeed, in this work, we use simplified analytic Keplerian orbits, which allow us to assume a constant transfer function for the duration of the measurement across the frequency range.
Note that we also have reported Bayes factors computed from a single frequency-domain realization of the instrument noise and SGWB signal, with spectra corresponding to their PSDs. This should constitute a good estimate of the expectation value of the Bayes factors, taken over the noise and signal realizations. Finally, we have assumed conservative uncertainties in the spectral estimates, as we adopt a lower bound for the effective number of degrees of freedom in the Wishart likelihood. This can be refined in the future by accurately including the effect of windowing.
The pipeline presented in this work is versatile and can fit various types of models while adapting their dimensionality. In addition, more recent version of the software can handle time-varying stochastic components, making it suitable for inferring both the non-stationary properties of the instrumental noise as well as anisotropic SGWB signals. A prime example is the stochastic signal generated by the unresolved Compact Galactic Binaries. We believe that such a tool will be useful for the analysis of the future data of the LISA mission.
@@ Acknowledgments
We acknowledge useful and inspiring discussions with F. Pozzoli, R. Buscicchio, M. Muratore, N. Dam Quang, O. Hartwig, M pieroni, A. Santini, R. Meyer, S. Babak, C. Caprini, N. Cornish, H. Inchauspé, A. Petiteau, M. Besançon, and G. Nardini and the members of the LISA Distributed Data Processing Center (DDPC). NK was supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the 4th Call for HFRI Research Projects to support Post-doctoral Researchers (Project Number: 28418).
@@
For the analysis and modelling of the LISA constellation, we have used the backgrounds software [^1] The package is hosted on a private instance of Gitlab, at https://gitlab.in2p3.fr/qbaghi/backgrounds.. We have also utilized the CudAkima package by A. Santini [44], and the Eryn sampler for its RJ-MCMC capabilities [^2] Publicly available at https://github.com/mikekatz04/Eryn..
@@ Appendix A Complementary figures for GW parameter posteriors
In this section, we present complementary figures that are relevant to the analysis performed for the production of figure 2. As already mentioned in section 3, we perform parameter estimation analyses for different injections of power-law SGWB signals, under different prior assumptions on the instrumental noise, while using analytic models and flexible models for the signal and noise components of the data. As an example, in figure A1, we show the parameter estimation results for a stochastic signal with $\Omega_{0}=5\times 10^{-15}$ and $n=2$. By the result shown in figure 2, we should expect this signal to be detectable with very high confidence when using analytic models for all components (red data of figure A1), while the detectability should drop when using the more flexible spline models for the PSD of the noise components (blue data of same figure). The two panels of figure A1 demonstrate the effect of the prior choice.
@@
In figure A2, we show the resulting combined posteriors on the spline knots positions and amplitudes for each model with $k$ knots, as sampled with trans-dimensional MCMC methods. The left panel corresponds to the TM noise component, while the right panel to the OMS noise component. We remind the reader that all noises are set to be equal across the three spacecraft, and therefore are being fit simultaneously with the same model in our analysis. The shaded gray area represents the wide uniform prior around the injected signal. Since we fit for deviations from the injected spectra, there is no need for high-order models to capture the complexity of the data, and therefore the number of spline knots estimated converges to the lower prior bound of $k=2$. This is shown in the embedded histograms in both panels of figure A2.
In figure A3 we demonstrate how the power induced by a relatively weak isotropic SGWB signal can be fully absorbed by the flexible noise models. In particular, we inject a marginally detectable signal with $\Omega_{\mathrm{GW}}=10^{-13},\,n=0$, and then we analyze the data with flexible spline models for the noise. For this investigation we only fit for the noise. In the left panel A3a we show the noise budget as estimated with the flexible noise models, while in the right panel A3b we focus on the deviations of the estimated spectra from the injected true noise. It is quite obvious that part of the signal power has been absorbed by the model which aims to characterize the TM noise contribution. This effect is naturally more prevalent for weak stochastic signals, and for uninformative uniform priors for the noise power spectra.
@@
Figure A1: Comparison of the parameter estimates of the SGWB signal ($\Omega_{0}=5\times 10^{-15},\,n=2$) with a power-law model, under different prior assumptions, and with different models for the noise (parametric and flexible, based on spline interpolation). (a) Using the informative prior based on the $\mathcal{SN}$ distribution, and (b) using the flat uninformative prior. See text for details.
@@
Figure A2: Examples of the combined posterior of the spline knots amplitudes $\log S_{j,k}$ and position $\log f_{j,k}$ for each $j$ spline number and given model order $k$, for the two instrumental noises, (a) for the TM and (b) the OMS noises. The shaded area represents the uniform prior used for this particular run, while the dashed black line denotes the injected true spectrum. The embedded figure represents the estimated model order, or the number of spline knots, as estimated via RJ-MCMC sampling. This analysis corresponds to a signal injection with $\Omega_{0}=5\times 10^{-15},\,n=2$ and the final signal estimate shown in figure A1b. See text for more details.
Figure A3: Demonstration of a case where the power induced by a relatively weak isotropic SGWB signal is be fully absorbed by the flexible noise models. For this investigation we inject a SGWB signal with $\Omega_{\mathrm{GW}}=10^{-13},\,n=0$, while we fit for only the noise using flexible spline models. (a) The reconstructed noises of the TDI X channel are shown with the green heatmap for the TM and with the pink for the OMS noises respectively. The injected true spectra are shown with the dashed and dot-dashed gray lines respectively, while the injected SGWB signal with red. (b) A zoom in on the deviations from the injected noise spectra. The gray dotted line represents the baseline around zero. The dashed red line represents the sum of the true injected spectra of TM and SGWB signal. It is evident that the power induced by the SGWB signal is absorbed by the TM noise model.